OpenStudy (anonymous):

Consider the following circle. (If an answer does not exist, enter DNE.) x2 + y2 − 10x + 4y + 15 = 0 (a) Find the x-intercepts. (b) Find the y-intercept.

OpenStudy (anonymous):

Solve the two quadratic equations, one with x=0, the other with y=0.

OpenStudy (anonymous):

To find the x-intercept, plug 0 into y and solve for x. To find the y-intercept, plug 0 into x and solve for y.

OpenStudy (anonymous):

Found those solutions yet. @tymiller ?

OpenStudy (anonymous):

i put them into the equations and my online hw ste said the answers were wrong

OpenStudy (anonymous):

trying it again

OpenStudy (anonymous):

What solutions did you come up with?

OpenStudy (anonymous):

(x^2-10x-25)+(y^2+4y-4)=radical 14 x=10+/- radical 10^2-4(1)(-25)/2 x=5+/-5 radical 2 y=-4+/-radical 4^2-4(1)(-4)/2

zepdrix (zepdrix):

Hmm

OpenStudy (anonymous):

y=-2+/-4 radical 2

zepdrix (zepdrix):

(x-5)^2 + (y+2)^2 = 14 Is this the equation of the circle you were able to get it to?

OpenStudy (anonymous):

Yes, hmmm

zepdrix (zepdrix):

Because I see that you were able to complete the square on x and y, after you do that, write it that way :D

zepdrix (zepdrix):

the x term should have a +25, not -25 to complete that square :D minor detail

OpenStudy (anonymous):

From \[x^2 + y^2 − 10x + 4y + 15 = 0\] Y-intercepts are where: \[y^2 + 4y + 15 = 0\] X-intercepts are where: \[x^2− 10x + 15 = 0\] Neither are factorable, so use quadratic formula.

OpenStudy (anonymous):

Completing the square isn't necessary, but can be useful. It would make it easier to graph, which can help visualize the solutions.

OpenStudy (anonymous):

still wont except the answer.

OpenStudy (anonymous):

From completing the square, you can see that the circle has center at (5,-2) with a radius of √14. The graph looks like this: 

OpenStudy (anonymous):

yea i got that part

OpenStudy (anonymous):

From the graph, you see that you'll get two solutions for the x-intercepts, and no real solutions for the y-intercepts. (again, completing the square and graphing is not necessary; use quadratic formula on the two equations for x and y.)

OpenStudy (anonymous):

:( Webassign still doesnt like the answers

zepdrix (zepdrix):

Hmm what did you put in? :)

zepdrix (zepdrix):

\[x=\sqrt14+5\]\[x=-\sqrt14+5\] Are you getting these for solutions? That's what I came up with :d lemme know if yer still confused.

OpenStudy (anonymous):

those answers do not work i have tried

zepdrix (zepdrix):

Hmmmmm